OpenStudy (anonymous):

find the slop of the line 3x+5y=-3 write your answer in a simplest form

OpenStudy (anonymous):

okay, so in order to solve this problem you need to switch its form to y=mx+b form. so in order to do that you need to move the 3x to the other side using subtraction and then divide everything by 5 so it would look like this after words... \[y=-\frac{ 3 }{ 5 }x-\frac{ 3 }{ 5 }\] in y=mx+b form M is your slope so the slope is \[-\frac{ 3 }{ 5 }\]